OpenStudy (anonymous):

a car travels in a straight line with vatying velocity given by v={t - t0}m/s for 0<_<_2t0 where t=time and t0= time at which velocity is zero the distance travelled by the car in this interval is?

OpenStudy (anonymous):

option are t0^2 2t0^2 3/2t0^2 0

OpenStudy (inkyvoyd):

Not sure. Can we go back to the last qusetion?

OpenStudy (anonymous):

the answer is a

OpenStudy (inkyvoyd):

I'm not sure how to answer this question. @dumbcow ?

OpenStudy (anonymous):

i think we have to use vector

OpenStudy (anonymous):

x= \[x=x0+v0 \times t + at ^{2}/2\] since x0 = 0 - there is no any notion on initial position of object v0 = v(t0)=0 it is given in problem we get \[x=at ^{2}/2\]

OpenStudy (anonymous):

let's find acceleration

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

at=v => a=v/t v= t-t0 a=(t-t0)/ t

OpenStudy (dumbcow):

i guess i would go with \[d = \int\limits_{0}^{2t_0}v(t) dt\]

OpenStudy (anonymous):

Now try to put expression for a and t= 2t0 in \[x=at ^{2}/2\]

OpenStudy (anonymous):

i get it thanxx to both of u another question plzz view it